OpenStudy (anonymous):

how do i compute the exact value of sin(-10pi/3)

OpenStudy (jamesj):

like I pointed out to earlier sin(x) = sin(x + 2pi) So sin(-10pi/3) = sin(-10pi/3 + 4pi) = sin(2pi/3) Now figure out that.

OpenStudy (mathmate):

sin(-10pi/3) =sin(-10pi/3+4pi) (periodic function) =sin(2pi/3) =sin(pi/3) =sqrt(3)/2